Grade 12th passTrigonometry

When sin9a/cos27a+sin3a/cos9a+sina/cos3a=k(tan27a-tana) is defined,then k =

Krishnaveni
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
(sinA / cos3A) 
= (1/2) * [(2 * sinA * cosA) / (cosA * cos3A)] 
= (1/2) * [(sin2A) / (cosA * cos3A)] 
= (1/2) * [sin(3A - A) / (cosA * cos3A)] 
= (1/2) * [(sin3A * cosA - cos3A * sinA) / (cosA * cos3A)] 
= (1/2) * [{(sin3A * cosA) / (cosA * cos3A)} - {(cos3A * sinA) / (cosA * cos3A)}] 
= (1/2) * [(sin3A/cos3A) - (sinA/cosA)] 
= (1/2) * (tan3A - tanA) 

so, (sinA / cos3A) = (1/2) * (tan3A - tanA) ---------(1) 

Now putting 3A in place of A in equation (1) we get 
(sin3A / cos9A) = (1/2) * (tan9A - tan3A) ------------(2) 

Again putting 9A in place of A in equation (1) we get 
(sin9A / cos27A) = (1/2) * (tan27A - tan9A) ------------(3) 

Now doing (1)+(2)+(3) we get 
(sinA / cos3A) + (sin3A / cos9A) + (sin9A / cos27A) 
= (1/2) * (tan3A - tanA + tan9A - tan3A + tan27A - tan9A) 
= (1/2) * (tan27A - tanA) 
 
Regards
Arun